### v1

# Cost-Efficient Subjective Task Annotation and Modeling through Few-Shot Annotator Adaptation

Abstract
In subjective NLP tasks, where a single ground truth does not exist, the inclusion of diverse annotators becomes crucial as their unique perspectives significantly influence the annotations. In realistic scenarios, the annotation budget often becomes the main determinant of the number of perspectives (i.e., annotators) included in the data and subsequent modeling. We introduce a novel framework for annotation collection and modeling in subjective tasks that aims to minimize the annotation budget while maximizing the predictive performance for each annotator. Our framework has a two-stage design: first, we rely on a small set of annotators to build a multitask model, and second, we augment the model for a new perspective by strategically annotating a few samples per annotator. To test our framework at scale, we introduce and release a unique dataset, Moral Foundations Subjective Corpus, of 2000 Reddit posts annotated by 24 annotators for moral sentiment. We demonstrate that our framework surpasses the previous SOTA in capturing the annotators’ individual perspectives with as little as 25% of the original annotation budget on two datasets. Furthermore, our framework results in more equitable models, reducing the performance disparity among annotators.
1 Introduction
The common pipeline for supervised learning in Natural Language Processing (NLP) starts by collecting annotations from multiple annotators. These annotations are often aggregated through majority voting (Talat and Hovy, 2016) to construct a ground truth or gold standard on which the subsequent modeling is performed. In recent years, researchers have advocated for a transition from single ground-truth labels to annotator-level modeling, aiming to capture diverse perspectives, enhance contextual understanding, and incorporate cultural nuances (Uma et al., 2021), and have proposed different frameworks that take into account unique perspectives of the annotators by modeling them as separate subtasks (Davani et al., 2022; Kanclerz et al., 2022).
The impact of individual annotators’ backgrounds and life experiences on annotations in subjective tasks signifies the importance of incorporating a diverse set of annotators. Nevertheless, the primary constraint on achieving this diversity is often the annotation budget, limiting the number and, consequently, the diversity of perspectives considered. In this paper, we introduce a novel framework for annotation collection and modeling in subjective tasks. Our framework is designed to minimize the annotation budget required to model a fixed number of annotators, while maximizing the predictive performance for each annotator.
Our framework operates in two stages. In the first stage, data is collected from a small pool of annotators. This data serves as a foundation for building a multitask model that captures the general patterns for the task and provides a signal of differences among individual annotators. Informed by the first stage annotations, the second stage involves collecting a few samples from each new annotator that best capture their differences from the general patterns. We use this data to augment the model from the first stage to learn the new annotators’ perspective from a few examples (Figure 1).
We introduce a unique dataset that enables the study of detecting moral content, an understudied subjective task, at a scale that was not possible before [^1] The dataset will be released as part of the accepted paper. The Moral Foundations Subjective Corpus (MFSC) is a collection of 2000 Reddit posts, each annotated by 24 annotators for moral content along with annotators’ responses to a range of psychological questionnaires (§4.1).
We use MFSC in conjunction with the Brexit Hate Dataset (Akhtar et al., 2021) to extensively study each component of our proposed framework. First, we empirically investigate the effect of pooling data from varying numbers of annotators in a multitask model (§5.1) and demonstrate that increasing the number of annotators does not improve annotator-level modeling in a multitask model. Second, we showcase the efficacy of our framework in capturing diverse annotator perspectives under budget constraints (§5.2). Third, in §5.3, we study the impact of various sample selection strategies. Our framework achieves a 4.3% increase in $F_{1}$ score with access to just 25% of the annotation budget in moral sentiment prediction. Furthermore, our results demonstrate a 6.7% improvement over the previous state-of-the-art in hate speech detection when using only 50% of the original annotation budget. Next, we show that our proposed framework consistently yields a more equitable model by minimizing the performance disparity across various annotators (§5.4). Specifically, on the lowest budget scenarios, our approach reduces the standard deviation in performance across annotators by 7.5% and 1.1% on hate speech detection and moral foundation classification, respectively. Finally, we extend our analysis to investigate whether the selection of the initial set of annotators in the first stage of our framework is related to the model’s performance (§5.5).
Our experiments on two subjective datasets revealed that our framework consistently surpasses previous state-of-the-art models with access to as little as 25% of the original annotation budget. In addition, our framework produced more equitable models with reduced performance disparities among the annotators. By minimizing data requirements, our cost-efficient framework for subjective tasks enables us to scale the number of included annotators and, hence, improve the diversity of captured perspectives. Furthermore, the two-stage design of our framework facilitates the integration of new annotators into pre-existing datasets.
2 Related Work
Subjective Tasks in NLP: In recent years, the variety of tasks for which NLP is used has significantly expanded. In many of these tasks, a single ground truth does not exist, making them inherently subjective in nature. In subjective tasks, researchers have argued that disagreements in particular labels should not be treated as statistical noise (Larimore et al., 2021; Pavlick and Kwiatkowski, 2019; Plank, 2022), as they are often indicative of individual differences which are driven by different backgrounds and lived experiences of the annotators (Akhtar et al., 2019; Plank et al., 2014; Prabhakaran et al., 2021; Díaz et al., 2018; Garten et al., 2019; Ferracane et al., 2021). For example, Davani et al. (2023) revealed how the stereotypes of annotators influence their behavior when annotating hate speech. In a similar context, Sap et al. (2021) demonstrate that annotators’ identity and beliefs impact their ratings of toxicity. Sang and Stanton (2022) conducted a study showing that differences in age and personality among annotators result in variations in their annotations. Larimore et al. (2021) explored how annotators’ perceptions of racism differ based on their own racial identity. Basile (2020) calls for a paradigm shift away from majority aggregated ground truths, and towards representative frameworks preserving unique perspectives of the annotators. In their later work, Basile et al. (2021) define the phenomena of Data Perspectivism, and share recommendations and outlines to advance the perspectivist stance in machine learning.
Capturing the Perspectives: To capture annotator-level labels, Akhtar et al. (2020) proposed dividing annotators into groups based on similar personal characteristics and creating different sets of gold standards for each group. Kanclerz et al. (2022) and Deng et al. (2023) incorporated knowledge about annotators into their models to make them personalized. Davani et al. (2022) propose a multitask approach, modeling each annotators’ perspective as a subtask, while having a shared encoder across the subtasks. Baumler et al. (2023) and Wang and Plank (2023) propose active learning methods for reducing the budget of data collection by proposing methods for collecting samples based on model confidence and annotators’ disagreement. Casola et al. (2023) also proposes ensembling perspective-aware models based on their confidence.
3 Method
Figure 1: Left: The baseline approach for annotator-level modeling, in full and reduced budget scenarios. Right: Our two-stage proposed framework, designed to achieve the outlined objectives
Problem Formulation: To formalize the task, suppose we have a set of annotators $\mathcal{A}=\{a_{1},...,a_{n}\}$ and input texts $X=\{x_{1},x_{2},...,x_{m}\}$ and their corresponding annotations $Y=\{y_{1},y_{2},...,y_{m}\}$. Let $D=\{D_{a_{i}}|a_{i}\in\mathcal{A}\}$ be the entire annotations and $D_{a_{i}}=\{X_{a_{i}},Y_{a_{i}}\}$ denote data collected from annotator $a_{i}$. Then the budget $B=|D|$ is defined as the total number of annotations collected. Let =$\mathcal{F}=\{f_{a_{i}}|a_{i}\in\mathcal{A}\}$ and $f_{a_{i}}$ denote the model capturing labels assigned by annotator $a_{i}$.
Proposed Framework: We design our framework with two objectives: first, maximizing the average performance over all annotators. Second, minimizing the budget ($B$) required to achieve the first goal. The second objective allows us to increase the number of annotators’ perspectives ($|\mathcal{A}|$ ) captured with a given budget. Our framework design is based on two key intuitions. Firstly, as we show in Section 5.1, multitask learning, which has often been treated as the upper bound by previous work, does not improve in performance as the number of annotators grows. Secondly, even in subjective tasks, there exists a substantial number of texts on which annotators mostly agree, particularly when these texts are randomly drawn from a source. Therefore, obtaining many annotations on such instances is not beneficial in learning a new perspective. In line with these intuitions, our framework consists of two stages (Figure 1). In the first stage, we learn the commonalities between annotators through a multitask model $\mathcal{F}_{mtl}$. A crucial difference of our approach in comparison to previous multitask methods is that we only collect annotations from a small subset of annotators $\mathcal{A}_{mtl}\subset\mathcal{A}$. In the second stage, we learn the perspectives of new annotators $\mathcal{A}_{fs}=\mathcal{A}-\mathcal{A}_{mtl}$ with only a few shots. Specifically, we collect annotations for $k$ input texts $\mathcal{S}(X)\subset X$, where $\mathcal{S}$ is a sampling function that ideally helps in capturing patterns specific to individual annotators’ perspectives. Let $D^{fs}_{a_{i}}=\{(x,y_{ai})|x\in\mathcal{S}(X)\}$ and $|D^{fs}_{a_{i}}|=k<<|D_{a_{i}}|$. We initialize $\mathcal{F}_{\mathcal{A}_{fs}}$ with $\mathcal{F}_{mtl}$ and train it on $D^{fs}_{a_{i}}$.
Sampling Function ($\mathcal{S}$): We explore four different sampling functions: 1) $\mathcal{S}_{rand}$: selects a random sample for each annotator 2) $\mathcal{S}_{mv}$: selects a balanced sample determined by the majority vote of the annotators. For a set of annotators $\mathcal{A}_{mtl}$, we calculate the majority vote among these annotators and select $k$ samples that have an equal number of each label based on that majority vote. 3) $\mathcal{S}_{dis}$ selects the samples from $\mathcal{A}_{mtl}$ with highest disagreement score, and 4) $\mathcal{S}_{bal}$ acts as an oracle, selecting a balanced sample based on a specific annotator’s label, not the majority vote. Therefore, if we have a new annotator, $\mathcal{S}_{bal}$ would select a balanced sample based on the annotations of that specific annotator. One frequent challenge in some subjective tasks is the heavy imbalance in class frequencies. Hence, we chose $\mathcal{S}_{mv}$ and $\mathcal{S}_{bal}$ to provide a more balanced sample to the few-shot model for each annotator. We added $\mathcal{S}_{dis}$ with the goal of providing samples that differentiate the individual annotator perspectives to the model. We use the “item disagreement” and “annotator disagreement” measure from Davani et al. (2023) to select samples in $\mathcal{S}_{dis}$.
4 Experiments
4.1 Datasets
For a dataset to be well-suited for our experiments it should meet the following four constraints: 1) the dataset should pertain to a subjective topic, one that allows for different perspectives, 2) individual annotators’ labels should be available, 3) there should be enough annotators to allow experimentation with varying budgets, and 4) each sample should be annotated by all annotators. This will allow us to ensure any observed performance gains are due to the used methods and are not driven by the specific samples annotated by each annotator. Below, we describe the two datasets: the MFSC, which we collected specifically to enable research on subjective tasks, and the Brexit Hate Dataset.
Moral Foundations Subjective Corpus (MFSC): We introduce a new dataset, Moral Foundations Subjectivity Corpus (MFSC), consisting of 2000 Reddit posts annotated by 24 annotators for moral sentiment based on the Moral Foundations Theory (MFT; Graham et al., 2013; Atari et al., 2022). Morality is widely acknowledged to be a subjective concept, strongly influenced by cultural backgrounds Graham et al. (2016), one that has not been explored much in the NLP community. We asked annotators to label each text for the specific moral concern (i.e., Purity, Harm, Loyalty, Authority, Proportionality or Equality), and if one of the concerns is chosen, we set the label as moral, else as non-moral; This dataset was collected following the same procedure as Trager et al. (2022). This dataset also contains additional metadata information, such as confidence for each instance using a 3-level measure (i.e., confident, somewhat confident, and not confident). We also collected annotator responses for the “Big Five Inventory-2-Short” questionnaire (Soto and John, 2017). MFSC provides an opportunity to explore the subjective nature of morality. The substantial number of annotators in this dataset along with questionnaire responses enables future researchers to investigate the modeling of subjective tasks on a larger scale. The demographics of the annotators is shown in Appendix A.1. In our experiments, we use the annotations for the moral label. In addition, we evaluate our framework on Care label which is discussed in Appendix B.2.
Brexit Hate dataset: Hate speech detection has become one of the primary subjective tasks studied in the NLP community (Akhtar et al., 2019; Sang and Stanton, 2022; Sap et al., 2021). The Brexit Hate dataset (Brexit) proposed by Akhtar et al. (2021), consists of 1,120 English tweets collected with keywords related to immigration and Brexit. The dataset was annotated with hate speech (in particular xenophobia and islamophobia), aggressiveness, offensiveness, and stereotype, by six annotators belonging to two distinct groups: a target group of three Muslim immigrants in the UK, and a control group who were researchers with Western background. For our experiments, we use the overall hate label. Additional dataset statistics can be found in Appendix A.
4.2 Experiment Setup
We designed our experiments to study the impact of each component of the framework towards our two objectives: maximizing average performance and minimizing annotation budget.
We use multitask learning (MTL) on all the annotators as our baseline and assess the efficacy of our framework compared to this baseline in capturing individual annotators’ perspectives under a range of budget constraints. Specifically, for our approach, we vary the budget $B$ by changing the size of $|\mathcal{A}_{mtl}|$. Recall that $B=|D|=\sum|D_{a_{i}}|$ and $|D^{fs}_{a_{i}}|=k<<|D_{a_{i}}|$. Also, recall that under our proposed framework the annotators $\mathcal{A}$ are divided into two sets $\mathcal{A}_{mtl}$ and $\mathcal{A}_{fs}$. Since the cost of annotating a few samples per new annotator is negligible ($\frac{|D^{fs}_{a_{i}}|}{|D_{a_{i}}|}$ is close to 0) the budget under our proposed framework can be reduced to
$B_{ours}$ $\approx\sum_{a_{i}\in\mathcal{A}_{\text{mtl}}}|D_{a_{i}}|$
$=\frac{\sum_{a_{i}\in\mathcal{A}_{\text{mtl}}}|D_{a_{i}}|}{\sum_{a_{i}\in\mathcal{A}}|D_{a_{i}}|}\times B$
$=\frac{|\mathcal{A}_{mtl}|}{|\mathcal{A}|}\times B$
For example, the MFSC dataset has $|\mathcal{A}|=24$ annotators. Hence, $25\%B$ shows the scenarios where $|\mathcal{A}_{mtl}|=6$. Whereas, for the baseline, we vary the budget $B$ by changing the size of $D_{a_{i}}$ for all annotators. In the given example, a $25\%B$ for the baseline means using only $25\%$ of $D_{a_{i}}$ for each $a_{i}$.
To ensure that our results are not driven by the specific choices of $\mathcal{A}_{mtl}$, we run our experiments for each budget on multiple samples of $\mathcal{A}_{mtl}\subset\mathcal{A}$. Specifically, we run our models with all possible choices of $\mathcal{A}_{mtl}$ for Brexit dataset and 20 different samples of $\mathcal{A}_{mtl}$ for the MFSC dataset.
For each annotator $a_{i}$, $F^{a_{i}}_{1}$ denotes the performance on predicting $a_{i}$’s labels. We use $F^{fs}_{1}$ and $F^{mtl}_{1}$ to denote the average of $F^{a_{i}}_{1}$ scores when $a_{i}\in\mathcal{A}_{fs}$ and $a_{i}\in\mathcal{A}_{mtl}$ respectively. For our framework, we also calculate the overall performance for all annotators $F^{\text{overall}}_{1}$ as the weighted average of $F^{fs}_{1}$ and $F^{mtl}_{1}$.
4.3 Implementation Details
We employ RoBERTa-base as the base model for our methods (Liu et al., 2019). All multitask models undergo hyperparameter tuning, specifically for learning rate and weight decay, and are trained for 5 epochs. We employ a weighted random sampler for the Brexit dataset to account for the imbalance in the labels of each annotator. Our training batch size is 128, and we use AdamW optimizer. The best hyperparameters of the best-performing model are then applied in the subsequent fine-tuning step. Few-shot models are trained for 50 epochs, during which all model parameters are updated. All models’ training is repeated with three random seeds. For the Brexit dataset, we adopt the predefined train, validation, and test splits provided within the dataset  [^2] https://le-wi-di.github.io/. For the MFSC dataset, we allocate 80% for training, 10% for validation, and the remaining 10% for testing.
5 Results and Analysis
5.1 Few Annotators are Enough
$metric=F_{1}^{overall}\uparrow$ Brexit MFSC
$50\%$ $66\%$ $83\%$ $100\%$ $25\%$ $50\%$ $75\%$ $100\%$
$X\%\times\lvert D_{a_{i}}\rvert$ $50\%\lvert D_{a_{i}}\rvert$ $66\%\lvert D_{a_{i}}\rvert$ $83\%\lvert D_{a_{i}}\rvert$ $\lvert D_{a_{i}}\rvert$ $25\%\lvert D_{a_{i}}\rvert$ $50\%\lvert D_{a_{i}}\rvert$ $75\%\lvert D_{a_{i}}\rvert$ $\lvert D_{a_{i}}\rvert$
MTL 0.417(0.049) 0.449(0.027) 0.418(0.018) 0.431(0.014) 0.763(0.016) 0.773(0.011) 0.772(0.015) 0.776(0.004)
$X\%\times\lvert\mathcal{A}\rvert$ $50\%\lvert\mathcal{A}\rvert$ $66\%\lvert\mathcal{A}\rvert$ $83\%\lvert\mathcal{A}\rvert$ $25\%\lvert\mathcal{A}\rvert$ $50\%\lvert\mathcal{A}\rvert$ $75\%\lvert\mathcal{A}\rvert$
Ours $k=16$ 0.422(0.012) 0.44(0.025) 0.455(0.007) 0.795(0.009) 0.79(0.006) 0.785(0.005)
Ours $k=32$ 0.428(0.006) 0.447(0.019) 0.452(0.016) 0.795(0.01) 0.791(0.006) 0.785(0.004)
Ours $k=64$ 0.433(0.012) 0.451(0.015) 0.456(0.013) 0.797(0.009) 0.791(0.006) 0.785(0.004)
Ours $k=128$ 0.439(0.015) 0.457(0.012) 0.455(0.011) 0.798(0.008) 0.791(0.005) 0.786(0.004)
Table 1: Mean and standard deviation of $F_{1}$ scores on Brexit and MFSC dataset across three runs, with varying annotation budgets (%$\%B$). The budget allocation differs in the baseline and our framework; For example, $50\%\lvert D_{a_{i}}\rvert$ indicates selecting 50% of each annotator’s data, while $50\%\lvert\mathcal{A}\rvert$ denotes selecting $50\%$ of annotators and using all their data (i.e., 3 out of 6 annotators for Brexit, and 12 out of 24 annotators for MFSC).
Previous state-of-the-art methods for subjective tasks achieve performance gains by jointly learning annotations from multiple annotators, as opposed to independently learning each annotator’s perspectives Davani et al. (2022). In this analysis, we further explore whether adding more annotators to the multitask model results in a better model for all annotators. To answer this question, we look at the multitask performance $F^{mtl}_{1}$ as the number of annotators in the multitask model, $|\mathcal{A}_{mtl}|$, grows. There is no statistically significant correlation between the number of annotators and $F_{1}^{mtl}$ ($r=-0.28,p>0.05$). Figure 2 shows that average $F^{mtl}_{1}$ for all annotators does not increase as the number of annotators increases. Note that this pattern is consistent in both datasets and as Figure 2 shows, the $F^{mtl}_{1}$ scores are reliable with low standard deviations across models with different numbers of annotators. This observation motivates our first design choice; in the first stage of our framework, we can only rely on a small number of annotators to get a reliable multitask model.
Figure 2: Multitask models $F_{1}$ performance ($F_{1}^{mtl}$) as the number of annotators increase. Mean and standard deviation are reported across three repeated runs
5.2 Towards Better Performance with Less Annotation Budget
A successful framework for modeling subjective tasks should, above all, demonstrate the ability to accurately predict labels from all of the annotators. Therefore, we assess the overall performance ($F^{\text{overall}}_{1}$) of our proposed framework in comparison to the baseline of multitask learning. Recall the second component of our framework is to augment the multitask learning model of the first stage $\mathcal{F}_{mtl}$ to predict new annotators’ labels using only a few samples. To isolate the impact of the few-shot learning model from the specific choice of samples, we sample the $k$ shots randomly.
As shown in Table 1, our framework outperforms the multitask baseline on both datasets when it comes to predicting labels of all annotators. Specifically, on MFSC, our approach outperforms multitask learning using 100% of data with access to as little as 25% of the original annotation budget. On Brexit, our framework matches the performance of multitask learning on 100% budget with access to only 50% of the original budget. The findings demonstrate the success of our framework in achieving its dual objectives: enhancing performance across all annotators while reducing annotation budget requirements. Consequently, our framework facilitates increased diversity by incorporating additional annotator perspectives while adhering to a specific budget constraint. We also conduct an ablation study by omitting the first MTL stage and employing random few-shot sampling for each annotator (Appendix B.1).
$metric=F_{1}^{fs}\uparrow$ Brexit MFSC
$50\%$ $66\%$ $83\%$ $100\%$ $25\%$ $50\%$ $75\%$ $100\%$
$X\%\times\lvert D_{a_{i}}\rvert$ $50\%\lvert D_{a_{i}}\rvert$ $66\%\lvert D_{a_{i}}\rvert$ $83\%\lvert D_{a_{i}}\rvert$ $\lvert D_{a_{i}}\rvert$ $25\%\lvert D_{a_{i}}\rvert$ $50\%\lvert D_{a_{i}}\rvert$ $75\%\lvert D_{a_{i}}\rvert$ $\lvert D_{a_{i}}\rvert$
MTL 0.417(0.049) 0.449(0.027) 0.418(0.018) 0.431(0.014) 0.763(0.016) 0.773(0.011) 0.772(0.015) 0.776(0.004)
$X\%\times\lvert\mathcal{A}\rvert$ $50\%\lvert\mathcal{A}\rvert$ $66\%\lvert\mathcal{A}\rvert$ $83\%\lvert\mathcal{A}\rvert$ $25\%\lvert\mathcal{A}\rvert$ $50\%\lvert\mathcal{A}\rvert$ $75\%\lvert\mathcal{A}\rvert$
$k=16$ $\mathcal{S}_{bal}$ 0.442(0.003) 0.457(0.016) 0.458(0.044) 0.785(0.006) 0.781(0.004) 0.778(0.01)
$\mathcal{S}_{dis}$ 0.399(0.013) 0.412(0.006) 0.457(0.036) 0.793(0.01) 0.795(0.004) 0.785(0.01)
$\mathcal{S}_{mv}$ 0.409(0.018) 0.416(0.026) 0.442(0.055) 0.799(0.003) 0.794(0.004) 0.786(0.015)
$\mathcal{S}_{rand}$ 0.4(0.017) 0.414(0.045) 0.44(0.037) 0.815(0.026) 0.814(0.033) 0.805(0.033)
$k=32$ $\mathcal{S}_{bal}$ 0.454(0.013) 0.469(0.004) 0.461(0.026) 0.787(0.006) 0.782(0.003) 0.779(0.012)
$\mathcal{S}_{dis}$ 0.402(0.008) 0.413(0.02) 0.455(0.03) 0.796(0.01) 0.797(0.004) 0.787(0.011)
$\mathcal{S}_{mv}$ 0.404(0.029) 0.426(0.029) 0.461(0.024) 0.801(0.003) 0.795(0.003) 0.787(0.017)
$\mathcal{S}_{rand}$ 0.412(0.006) 0.436(0.026) 0.422(0.027) 0.816(0.028) 0.815(0.035) 0.807(0.031)
$k=64$ $\mathcal{S}_{bal}$ 0.462(0.01) 0.47(0.018) 0.467(0.011) 0.788(0.005) 0.784(0.001) 0.782(0.011)
$\mathcal{S}_{dis}$ 0.429(0.018) 0.458(0.018) 0.518(0.017) 0.797(0.011) 0.799(0.005) 0.788(0.012)
$\mathcal{S}_{mv}$ 0.409(0.017) 0.41(0.05) 0.467(0.011) 0.802(0.004) 0.797(0.002) 0.79(0.018)
$\mathcal{S}_{rand}$ 0.423(0.02) 0.447(0.016) 0.447(0.019) 0.818(0.027) 0.815(0.033) 0.809(0.033)
$k=128$ $\mathcal{S}_{bal}$ 0.498(0.009) 0.517(0.025) 0.524(0.021) 0.79(0.005) 0.785(0.001) 0.785(0.01)
$\mathcal{S}_{dis}$ 0.456(0.011) 0.476(0.026) 0.511(0.023) 0.798(0.009) 0.799(0.003) 0.786(0.012)
$\mathcal{S}_{mv}$ 0.425(0.028) 0.428(0.035) 0.46(0.025) 0.805(0.004) 0.798(0.003) 0.79(0.019)
$\mathcal{S}_{rand}$ 0.433(0.023) 0.466(0.006) 0.443(0.027) 0.819(0.026) 0.816(0.032) 0.81(0.031)
Table 2: Few-shot $F_{1}$ results ($F_{1}^{fs}$) on BREXIT and MFSC datasets across varying percentages of the full budget (%$\%B$). Mean and standard deviation calculated over three repeated runs.
5.3 Few-shot Sampling Strategies
Recall that in the second stage of our framework, we only select a small subset of the input texts $\mathcal{S}(X)\subset X$ to be annotated by the new annotators (Figure 1). Intuitively, we are relying on the first stage of our framework (i.e., multitask learning) to learn the commonalities between annotators. Hence, the goal of the selection function $\mathcal{S}(X)$ is to choose input texts that best capture individual annotators’ differences. To study the impact of data selection strategies, we compare the fewshot performance ($F^{fs}_{1}$) of each strategy controlling for the number of shots $k$ (see §3 for definitions of $\mathcal{S(.)}$).
On Brexit dataset, the $\mathcal{S}_{bal}$ strategy, which selects a balanced sample for each annotator outperforms other approaches consistently with 50% budget. With 66% budget, $\mathcal{S}_{bal}$ strategy performs the best for all values of $k$. Notably, with this budget, the performance gap is closed by the $\mathcal{S}_{dis}$ for higher number of shots. Although the differences in performance are less pronounced for higher budget scenarios (i.e., the 83%$B$), $\mathcal{S}_{bal}$ and $\mathcal{S}_{dis}$ still achieve better results especially when more shots are available. We acknowledge that the $\mathcal{S}_{bal}$ strategy relies on access to a balanced sample for each annotator which is often not readily available. However, this study showcases the potential of this approach and future work can explore methods that attempt to select instances that approximate a balanced sample per annotator.
On the MFSC dataset, sampling strategies persistently outperform the multitask learning baseline in all scenarios. While different strategies seem to perform comparably, a crucial differentiating factor is the impact of various strategies on the stability of the results. Specifically, compared to $\mathcal{S}_{rand}$, all sampling functions result in more stable performances (i.e., lower standard deviation) across all budgets and values of $k$.
Overall, our results on two datasets show that the fewshot stage of our framework, results in models that consistently outperform the multitask learning baseline. Our experiments clearly demonstrate the importance of sample selection strategies both in terms of performance improvement and stability. Furthermore, our results on both datasets motivate the exploration of sampling strategies that can approximate a balanced sample per annotator.
5.4 Reduced Performance Disparities across Annotators
Ensuring a comprehensive representation of annotators’ viewpoints is crucial in modeling subjective tasks. To achieve this goal, a critical criterion is to create models that not only improve the aggregated performance but also demonstrate fair and equitable performance across all annotators. For example, if the $F_{1}$ scores of one model for two annotators are 0.6 and 0.8, respectively, while the second model scores 0.7 for both annotators, the latter is considered a better model. Although the average performance is the same for both models, the first model has a disparate negative impact on the first annotator.
Figure 3: Comparison of Annotator level F1 scores ($F_{1}^{a_{i}}$) on the Brexit dataset between MTL model and our framework, leveraging the $\mathcal{S}_{bal}$ sampling method for all budgets and shots
This is important because performance disparities among social groups (in our case annotators) can lead to biased models, limiting the system’s ability to accurately reflect diverse perspectives and potentially perpetuating inequalities in the outputs of subjective tasks (Buolamwini and Gebru, 2018). Merely relying on aggregated performance measures, like the average across all annotators, falls short of providing a complete understanding of how well the model is capturing annotators’ varying perspectives. For example, it is not clear whether the average performance is improving because an approach improves on capturing only a subset of annotators or for everyone. Hence, we look into the standard deviation of performance across all annotators $d=\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}(F^{a_{i}}_{1}-\overline{F^{overall}_{1}})^{2}}$. Lower standard deviations are indicative of more equitable models. As shown in Table 3, regardless of the sampling strategy, our approach results in lower performance variances compared to the MTL baseline on lower budgets (50% and 66%). Comparing the different sampling strategies on Brexit dataset, we observe that $S_{bal}$ achieves the most equitable models across all values of $k$. On MFSC, using a balanced sampling strategy consistently reduces the disparities among annotators compared to the MTL baseline for all budgets. Figure 3 visualizes the performance of our framework in comparison to the multitask learning baseline for each annotator. Notably, our framework improves the performance for the annotators in the non-Western control group (i.e., the first three annotators) while maintaining the performance of the rest of the annotators.
Overall, these results suggest that our proposed framework not only improves the overall performance of all annotators but also yields models that are more fair and equitable. Interestingly, we observe similar patterns in fairness and overall performance improvements. In other words, the configurations with the best overall performance (see section 5.3) are also the ones that have the least performance disparities among all annotators.
5.5 Annotator level analysis
Here we delve into the relationship between annotator-level variables. Recall that our framework is trained on $\mathcal{A}_{mtl}$ in the initial stage, followed by fine-tuning for each $a\in\mathcal{A}_{fs}$. Hence, a practical question arises: does the choice of the set $\mathcal{A}_{mtl}$ matter? In other words, would the similarity or divergence in perspectives among annotators in this set impact the performance on $\mathcal{A}_{fs}$? Investigating this is crucial, as identifying such an effect would necessitate a thoughtful selection of $\mathcal{A}_{mtl}$. To examine this, we conduct the following analysis:
  $metric=d\downarrow$ Brexit MFSC
$50\%$ $66\%$ $83\%$ $25\%$ $50\%$ $75\%$
  MTL .168 .139 .131 .128 .136 .127
$k=16$ $\mathcal{S}_{bal}$119 .127 .132 .119 .124 .124
$\mathcal{S}_{dis}$141 .134 .127 .129 .130 .126
$\mathcal{S}_{mv}$138 .137 .131 .131 .130 .127
$\mathcal{S}_{rand}$137 .139 .135 .137 .135 .130
$k=32$ $\mathcal{S}_{bal}$113 .125 .133 .120 .125 .124
$\mathcal{S}_{dis}$145 .138 .131 .131 .130 .126
$\mathcal{S}_{mv}$140 .137 .129 .130 .130 .127
$\mathcal{S}_{rand}$139 .135 .138 .136 .134 .129
$k=64$ $\mathcal{S}_{bal}$109 .119 .127 .119 .124 .123
$\mathcal{S}_{dis}$137 .126 .123 .131 .130 .126
$\mathcal{S}_{mv}$145 .146 .126 .130 .129 .127
$\mathcal{S}_{rand}$131 .128 .132 .136 .135 .129
$k=128$ $\mathcal{S}_{bal}$093 .108 .117 .117 .122 .121
$\mathcal{S}_{dis}$111 .120 .124 .130 .129 .127
$\mathcal{S}_{mv}$137 .142 .132 .126 .128 .127
$\mathcal{S}_{rand}$131 .127 .136 .134 .133 .128
[…]
Disagreement within $\mathcal{A}_{mtl}$ and performance on $\mathcal{A}_{fs}$: The aim of this analysis is to investigate whether there is a relationship between the disagreement within annotators in $\mathcal{A}_{mtl}$ and the performance of the newly adopted annotators in $A_{fs}$. To test this relationship, we employ a mixed-effects model to predict the performance of $a\in\mathcal{A}_{fs}$ by the agreement within $\mathcal{A}_{mtl}$ denoted as $d^{1}$ (Fleiss, 1971). The model controls for $k$, budget $B$, and agreement between $\mathcal{A}_{fs}$ and $\mathcal{A}_{mtl}$, denoted using $d^{2}$, incorporating random effects for $\mathcal{A}_{mtl}$ and $\mathcal{A}_{fs}$. The formula for this model is as follows:
$\begin{split}f_{ij}=&\beta_{0}+\beta_{1}d^{1}_{j}+\beta_{2}k_{ij}+\beta_{3}B_{j}\\ &+\beta_{4}d^{2}_{ij}+u_{0i}+v_{1j}+e_{ij}\end{split}$ (1)
where $f_{ij}$ denotes the performance of $i^{\text{th}}$ annotator in $A_{fs}$ on the model trained on a $j^{\text{th}}$ sample of $A_{mtl}$. The fixed effects coefficients are represented by $\beta_{0}$ to $\beta_{4}$, and the random effects for $i$ and $j$ are represented by $u_{0i}$, $v_{1j}$ respectively. $e_{ij}$ denotes the residual error term. To see the impact of sampling strategies, we run a total of four models, each corresponding to the performance results obtained from one of the strategies ($\mathcal{S}_{bal},\mathcal{S}_{dis},\mathcal{S}_{mv},\mathcal{S}_{rand}$).
The findings regarding Brexit indicate that there is no statistically significant effect of agreement within $\mathcal{A}_{mtl}$ ($d^{1}$) on the performance. For the MFSC dataset, a significant effect was observed only for results obtained from $\mathcal{S}_{bal}$ ($\beta_{1}=-0.052,SE=0.012,p<0.001$). This implies that a unit decrease in $d^{1}$, corresponding to moving from full agreement to full disagreement, is associated with a 0.052 increase in the $F_{1}$ score. This finding suggests that selecting a diverse $\mathcal{A}_{mtl}$ with high disagreement can potentially be advantageous.
6 Conclusion
We introduced a framework for annotation collection and annotator modeling in subjective tasks. Our framework aims to minimize the annotation budget required to model a fixed number of annotators while maximizing the predictive performance for each annotator. Our approach involves collecting annotations from an initial set of annotators and building a multitask model that captures general task patterns while signaling differences among individual annotators. Subsequently, we utilize the annotations from the first stage to select a small set of samples from new annotators that best highlight their deviations from the general patterns. Finally, we leverage this collected data to augment the initial model, enabling it to learn the new annotator’s perspective from a limited number of examples. We explored four distinct methods for few-shot sample selection and found that the most effective approach involves balanced sample selection. We introduced a new subjective task dataset Moral Foundations Subjective Corpus (MFSC ), of 2000 Reddit posts annotated by 24 annotators for moral sentiment which enabled us to test our framework in scale. Our experiments on MFSC and a hate speech dataset revealed that our framework consistently surpasses previous SOTA with access to as little as 25% of the original annotation budget. In addition, we showed that our framework yields more equitable models that reduce performance disparities among annotators. Our cost-effective framework for subjective tasks allows increasing the number of annotators, enhancing the diversity of perspectives captured, and facilitates the integration of new annotators into pre-existing datasets.
7 Limitations and Ethical Statement
We acknowledge that the datasets employed in our experiments are not representative of all annotator populations. While in MFSC we recruited a substantial number of annotators and efforts were made to diversify this pool, it is important to note that our sample is limited to undergraduate students at a private university in the US. Consequently, we advocate for the replication and extension of our work with non-student, non-US-based samples. Furthermore, we exclusively operate with English data and focus on datasets related to moral sentiment prediction and hate speech detection tasks. This may restrict the generalizability of our findings to a broader linguistic and thematic landscape. Despite these constraints, our research lays the groundwork for future research to extend and validate our approach across diverse languages and subjective NLP tasks. In our experiments, we do not consider the cost of collecting few-shot samples, as discussed in Section 4.2. We recognize that in certain cases, depending on the budget and the nature of the task, this assumption can be challenged. Even with the additional expense of annotating a few samples per new annotator, it is crucial to highlight that our proposed framework substantially reduces annotation cost, especially as the number of included perspectives grows.
In the MFSC dataset the annotators underwent four sessions of training, including guidance on avoiding potential adverse consequences of annotations, and were compensated at a rate of \$17 per hour. The study protocol received approval from the Institutional Review Board (IRB), and all annotators consented to both the terms outlined in an information sheet provided by the IRB about the study and the sharing of their responses to the psychological questionnaires along with their annotations. We emphasize that MFSC is created with the intention of exploring subjectivity and different perspectives in this context and it should not be used for any other purposes.
Appendix A Dataset details
Table 8 displays detailed information about each dataset, including examples and annotations of all annotators.
A.1 Demographics of MFSC Annotators
We aimed to diversify the annotators for MFSC dataset across gender, sexual orientation, religion, and race. Even though our dataset is not balanced across these dimensions, we strived to include representative annotators from a cross-section of the aforementioned demographics. The distribution of the annotators across the mentioned demographics is presented in Figure 4.
Figure 4: The abbreviations in the pie chart for race W stands for White or European American, B stands for Black or African American, H stands for Hispanic or Latino/Latinx, P stands for Native Hawaiian or Pacific Islander, A stands for Asian or Asian American, M stands for Middle Eastern or North African.
Appendix B Experiments and Results
B.1 Additional Baseline and Ablation Study
First, we conduct an ablation study where we omit the first stage MTL. Essentially, this model is equivalent to few-shot adaptation for each annotator initializing from a pre-trained model. The resulting $F_{1}$ scores are presented in Table 4. Comparing these scores with our framework using random few-shot sampling for $k=16$ from Table 2, we observe that our framework has a 21% gain for Brexit and a 15.5% gain for MFSC. This study demonstrates that the first stage MTL in our framework is crucial for its success.
In the second ablation study, we omit the second stage few-shot sample selection from our framework. In other words, in the second stage, we used the entire annotated samples for each annotator instead of selecting only a few samples. Note that this is equivalent to using 100% of the budget and serves as an upper bound to the performance achieved with an ideal sampling function. Additionally, we present Ensemble, a new baseline in which we train a separate model for each annotator using 100% of data from that annotator. Figure 5 presents a comparison of 3 different strategies, using 100% of the budget (MTL, Ensemble, and ours). On the Brexit Dataset (top) our framework has as much as 7.4% performance gain compared to Ensemble baseline (when using $\frac{4}{6}$ annotators in MTL), and for MFSC dataset our framework has as large as 5% gain compared to Ensemble baseline (when using $\frac{12}{24}$ of annotators in MTL). These results show that even considering the 100% budget our framework outperforms both baselines demonstrating the benefit of our two-stage design.
Figure 5: Comparision of our framework with baselines for the two datasets
Dataset $k=16$ $k=32$ $k=64$ $k=128$
Brexit 0.23 0.22 0.20 0.28
MFSC 0.66 0.73 0.71 0.77
Table 4: $F_{1}$ scores for few-shot baseline with random sampling $\mathcal{S}_{rand}$ when the first stage MTL is omitted.
B.2 Individual Moral Concern Results
We evaluate our framework on an additional binary label of Care moral concern from our MFSC dataset. This moral concern is defined as "Care/Harm: Intuitions about avoiding emotional and physical damage or harm to another individual. It underlies virtues of kindness, gentleness, and nurturing, and vices of meanness, violence, and abuse." (Trager et al., 2022). Table 5 presents the results for this task. Our framework outperforms the baseline MTL approach with $25\%$ and $50\%$ of the annotation budget. Notably, with only $25\%$ of the budget, our framework has a $1.4\%$ gain in $F_{1}$ score compared to MTL with 100% budget. The experiments were conducted with the same hyper-parameter tuning described in Section 4.3.
$metric=F_{1}^{Overall}\uparrow$ MFSC (Care)
$25\%$ $50\%$ $75\%$ $100\%$
$X\%\times\lvert D_{a_{i}}\rvert$ $25\%\lvert D_{a_{i}}\rvert$ $50\%\lvert D_{a_{i}}\rvert$ $75\%\lvert D_{a_{i}}\rvert$ $\lvert D_{a_{i}}\rvert$
MTL 0.474 0.476 0.49 0.469
$X\%\times\lvert\mathcal{A}\rvert$ $50\%\lvert\mathcal{A}\rvert$ $66\%\lvert\mathcal{A}\rvert$ $83\%\lvert\mathcal{A}\rvert$
$k=16$ $\mathcal{S}_{bal}$ 0.462 0.471 0.485
$\mathcal{S}_{dis}$ 0.46 0.467 0.485
$\mathcal{S}_{mv}$ 0.476 0.473 0.49
$\mathcal{S}_{rand}$ 0.469 0.468 0.482
$k=32$ $\mathcal{S}_{bal}$ 0.467 0.477 0.487
$\mathcal{S}_{dis}$ 0.463 0.463 0.483
$\mathcal{S}_{mv}$ 0.475 0.475 0.488
$\mathcal{S}_{rand}$ 0.47 0.468 0.484
$k=64$ $\mathcal{S}_{bal}$ 0.47 0.475 0.486
$\mathcal{S}_{dis}$ 0.467 0.471 0.478
$\mathcal{S}_{mv}$ 0.479 0.48 0.487
$\mathcal{S}_{rand}$ 0.472 0.477 0.49
$k=128$ $\mathcal{S}_{bal}$ 0.473 0.477 0.488
$\mathcal{S}_{dis}$ 0.474 0.474 0.481
$\mathcal{S}_{mv}$ 0.477 0.482 0.488
$\mathcal{S}_{rand}$ 0.483 0.481 0.487
Table 5: Overall $F_{1}$ scores on MFSC dataset, Care label, with varying annotation budgets (%$(\%B)$.
B.3 Overall Results of our Framework
$metric=F_{1}^{overall}\uparrow$ Brexit MFSC
$B=X\%\times\lvert\mathcal{A}\rvert$ $50\%\lvert\mathcal{A}\rvert$ $(\frac{3}{6})$ $66\%\lvert\mathcal{A}\rvert$ $(\frac{4}{6})$ $83\%\lvert\mathcal{A}\rvert$ $(\frac{5}{6})$ $25\%\lvert\mathcal{A}\rvert$ $(\frac{6}{24})$ $50\%\lvert\mathcal{A}\rvert$ $(\frac{12}{24})$ $75\%\lvert\mathcal{A}\rvert$ $(\frac{18}{24})$
$k=16$ $\mathcal{S}_{bal}$ 0.443(0.005) 0.454(0.015) 0.457(0.015) 0.777(0.002) 0.779(0.0) 0.78(0.003)
$\mathcal{S}_{dis}$ 0.421(0.01) 0.44(0.011) 0.457(0.008) 0.784(0.01) 0.787(0.004) 0.782(0.005)
$\mathcal{S}_{mv}$ 0.426(0.008) 0.441(0.019) 0.455(0.019) 0.789(0.004) 0.786(0.004) 0.783(0.006)
$\mathcal{S}_{rand}$ 0.422(0.012) 0.44(0.025) 0.455(0.007) 0.795(0.009) 0.79(0.006) 0.785(0.005)
$k=32$ $\mathcal{S}_{bal}$ 0.449(0.008) 0.458(0.009) 0.458(0.008) 0.779(0.002) 0.78(0.001) 0.78(0.003)
$\mathcal{S}_{dis}$ 0.423(0.008) 0.44(0.015) 0.457(0.016) 0.786(0.01) 0.788(0.004) 0.783(0.005)
$\mathcal{S}_{mv}$ 0.424(0.017) 0.444(0.02) 0.458(0.011) 0.791(0.004) 0.787(0.003) 0.783(0.007)
$\mathcal{S}_{rand}$ 0.428(0.006) 0.447(0.019) 0.452(0.016) 0.795(0.01) 0.791(0.006) 0.785(0.004)
$k=64$ $\mathcal{S}_{bal}$ 0.453(0.003) 0.458(0.016) 0.459(0.011) 0.78(0.003) 0.781(0.003) 0.781(0.004)
$\mathcal{S}_{dis}$ 0.436(0.01) 0.455(0.016) 0.468(0.01) 0.787(0.01) 0.789(0.004) 0.783(0.005)
$\mathcal{S}_{mv}$ 0.427(0.007) 0.439(0.026) 0.459(0.013) 0.791(0.005) 0.788(0.003) 0.784(0.007)
$\mathcal{S}_{rand}$ 0.433(0.012) 0.451(0.015) 0.456(0.013) 0.797(0.009) 0.791(0.006) 0.785(0.004)
$k=128$ $\mathcal{S}_{bal}$ 0.471(0.002) 0.474(0.018) 0.468(0.014) 0.781(0.002) 0.781(0.002) 0.782(0.003)
$\mathcal{S}_{dis}$ 0.45(0.008) 0.461(0.019) 0.466(0.016) 0.788(0.009) 0.789(0.003) 0.783(0.005)
$\mathcal{S}_{mv}$ 0.434(0.015) 0.445(0.022) 0.458(0.016) 0.793(0.005) 0.788(0.004) 0.784(0.007)
$\mathcal{S}_{rand}$ 0.439(0.015) 0.457(0.012) 0.455(0.011) 0.798(0.008) 0.791(0.005) 0.786(0.004)
Table 6: Overall $F_{1}$ results on Brexit and MFSC dataset for different budgets of annotation $(B)$, with various few shot sampling strategies; mean and standard deviation calculated over repeated runs.
For completeness of our analysis in Section 5.2, Table 6 presents the overall performance ($F^{\text{overall}}_{1}$) for all sampling strategies described in Section 3.
B.4 Effect of the Annotators’ Disagreement
In Figure 6 we demonstrate the impact of agreement (as a measure of similarity) between the first and second-stage annotators ($\mathcal{A}_{mtl}$ and $\mathcal{A}_{fs}$) on the performance of the model for the second stage annotators. Importantly, we do not observe performance degradation as the agreement between the two sets decreases.
B.5 Hardware Configuration
The experiments were conducted on 4 NVIDIA RTX A6000 gpus with 48GB RAM and overall they were completed in about one week.
(a) BREXIT
(b) MFSC
Figure 6: Each plot demonstrates the effect of a single annotator’s agreement with the initial set of annotators used for MTL training ($A_{mtl}$), on its $F_{1}$ score performance, when adopted as a few-shot task. The x-axis represents the agreement measure, and the y-axis represents the $F_{1}$ score. The darker color of the scatter plot corresponds to a higher number of positive labels provided by the respective annotator.
Appendix C Mathematical Symbols
Table 7 provides a directory of mathematical symbols used in our paper, along with their respective meanings, to facilitate ease of understanding for the reader.
Symbol Meaning
$\mathcal{A}_{fs}$ Annotators in MTL model
$\mathcal{A}_{mtl}$ Annotators adopted as few shot task
$\mathcal{S}_{mv}$ Sampling based on majority vote
$\mathcal{S}_{bal}$ Sampling based on balanced samples across classes
$\mathcal{S}_{dis}$ Sampling based on high disagreement of annotaions
$\mathcal{S}_{rand}$ Random sampling
$B$ Budget
$D$ All annotations for a dataset
$F^{fs}_{1}$ Avg. $F_{1}$ scores of the few-shot model for $\mathcal{A}_{fs}$
$F^{mtl}_{1}$ Avg. $F_{1}$ scores of the multi-task model for $\mathcal{A}_{mtl}$
Table 7: Mathematical notations used throughout the paper with their explanations
Dataset size #annotators Examples Annotations
Brexit 1120 6
"THE MAJORITY WILL NEVER allow the Mentally Ill Globalists to turn
the world into a SJW and Radical Islam SAFE SPACE #brexit #Trump2016"
"0,0,1,1,0,1"
"A muslim Mayor of London? What!? This PC Sickness has
become a pandemic. England turning into Little Asia. <url>"
"0,0,0,1,1,1"
Not all foreign people who wants to go to the uk have bad intentions.
Improve your law. The #Brexit isn’t gonna help your economy.
"0,0,0,0,0,0"
MFSC 2000 24
’m using the term *ethnicity* in all the meanings. A word doesn’t have an
inner, determined and true meaning; the meaning of a word is steadily
constructing by the speakers of a language. If people in the US use *ethnicity*
as a synonym of *race*, the term *ethnicity* acquires also that meaning,
moreover the anthropological meaning.I guess if they have French flags
celebrating Macron’s victory, they’d likely feel French.",
16 annotators : 0
8 annotators : 1
"More or less. Clinton was regarded as a ""female candidate"", but Le Pen
was a ""far-right candidate"". The left calls for diversity of
ethnicity/gender/etc, but conservatives don’t seem to count. ",
4 annotators : 0
20 annotators:1
"**ing bootlickers, all of them. They WANT a Trump monarchy.
All of them are traitors."
24 annotators: 1
Table 8: Statistics of the datasets used in our experiments, and examples from each alongside their annotations.